OBJECTIVE2 Exercise 2-31 Cost Classification Loring Company incurred the following costs last year: Sua Direct materials Factory rent Direct labor Factory utilities Supervision in the factory Indirect labor in the factory Depreciation on factory equipment Sales commissions Sales salaries Advertising Depreciation on the headquarters building Salary of the corporate receptionist Other administrative costs Salary of the factory receptionist S216,000 24,000 120,000 6,300 50,000 30,000 9,000 27,000 65,000 37,000 10,000 30,000 175,000 28,000 (Continued)
Required: 1. Classify each of the costs using the following table format. Be sure to total the amounts i each column. Example: Direct materials, $216,000. Product Cost Period Cost DirectDirect Manufacturing Selling Administrative Materials Labor Costs Overhead Expense Expense Direct materials $216,000 2. 3. 4. What was the total product cost for last year? What was the total period cost for last year? If 30,000 units were produced last year, what was the unit product cost?
OBJECTIVE 3 Exercise 2-36 Cost of Goods Sold Allyson Ashley makes jet skis. During the year, Allyson manufactured 94,00 jet skis. Finished goods inventory had the following units: January 1 December 31 6,800 7,200 Required 1. How many jet skis did Allyson sell during the year? 2. If each jet ski had a product cost of $2.200, what was the cost of goods sold last year? Use the following information for Exercises 2-37 and 2-38: In March, Chilton Company purchased materials costing $25,000 and incurred direct labor cost of $10,000. Overhead totaled $42,000 for the month. Information on inventories was as follows March 1 March 31 Materials Work in process Finished goods S14,000 8,000 9,000 $6,500 4,000 7,000
OBJECTIVE 3 Exercise 2-37 Direct Materials Used, Cost of Goods Manufactured Refer to the information for Chilton Company above. Required: 1. What was the cost of direct materials used in March? 2. What was the total manufacturing cost in March? 3. What was the cost of goods manufactured for March? OBJECTIVE 3 Exercise 2-38 Cost of Goods Sold Refer to the information for Chilton Company above. Required: What was the cost of goods sold for March?
Use the following information for Exercises 2-39 through 2-41. Jasper Company provided the following information for last year: Sales in units Selling price Direct materials Direct labor Manufacturing overhead Selling expense Administrative expense 280,000 S12 180,000 505,000 110,000 437,000 854,000 Last year, beginning and ending inventories of work in process and finished goods equaled an
Answer #1
Exercise 2-31: Product Cost Period Cost
Direct Materials Direct Labor Manufacturing Overhead Selling Expense Administrative Expense
Direct materials 216000
Factory rent 24000
Direct labor 120000
Factory utilities 6300
Supervision in factory 50000
Indirect labor in the factory 30000
Depreciation on factory equipment 9000
Sales commissions 27000
Sales salaries 65000
Advertising 37000
Depreciation on headquarters building 10000
Salary of the corporate receptionist 30000
Other administrative costs 175000
Salary of factory receptionist 28000
Total 216000 120000 147300 129000 215000
Total product cost for last year = 216000+120000+147300 = $            4,83,300
Total period cost = 129000+215000 = $            3,44,000
Unit product cost = 483300/30000 = $                  16.11
Exercise 2-36:
Number of jet skis sold = 6800+94000-7200 = 93600 Units
Cost of goods sold = 93600*$2200 = $    20,59,20,000
Exercise 2-37:
Cost of direct materials used = 14000+25000-6500 = $                32,500
Total manufacturing cost = 32500+10000+42000 = $                84,500
Cost of goods manufactured = 8000+84500-4000 = $                88,500
Exercise 2-38:
Cost of goods sold = 9000+88500-7000 = $                90,500
Note:
REQUIREMENT NOT GIVEN FOR JASPER COMPANY.